Can someone show me how to do this? The weak acid of vinegar, acetic acid, has a pKa value of 4.8. Determine the ratio of the acid form to the base form(i.e., determine [HA)/[A-]) at pH 6.0.
Answer #1

pH = 6.0

pH = 4.8

henderson hasselbalch equation

pH = pKa + log [A- / HA]

6.0 = 4.8 + log [A- / HA]

1.2 = log [A- / HA]

[A- / HA] = 10^1.2

            = 15.85

[HA /A-] = 1/15.85

              = 0.063

the ratio of the acid form to the base form = 0.063